OpenStudy (anonymous):

law of sines a distance BC of 160m is laid off on one side of the rier. found that B=109.3 C=15.8. find AB round to nearest meter

OpenStudy (jhannybean):

Interesting name* Hah.

OpenStudy (anonymous):

you or me?

OpenStudy (jhannybean):

You of course. Ok. analyzing this problem.

OpenStudy (anonymous):

well, i didn't know what to call myself so it's a song lyric i like actually "ooooo sometimes i get this feeling, yeah" kind of a cool older tune

OpenStudy (anonymous):

i thought it fit since math gives me a general funny feeling...all those unknowns (math joke)

OpenStudy (anonymous):

I used the law of sines but i keep coming up short on the answer, i really don't know how i could be messing this up unless i am needing to make another triangle than the one given

OpenStudy (danjs):

what is a rier?

OpenStudy (anonymous):

a river. sorry. slang for river maybe :)

OpenStudy (jhannybean):



OpenStudy (jhannybean):

Just making a guess at what the triangle might be.

OpenStudy (danjs):

oh

OpenStudy (anonymous):

54.9 for A

OpenStudy (danjs):

looks good, assuming he/she means angles B and C.. question is kinda ugly.. lol sorry

OpenStudy (jhannybean):

Law of sines: \[\frac{a}{\sin A} =\frac{c}{\sin C }\]

OpenStudy (anonymous):

i have some xtra credito word problems that are really ugly. do not make sense what so ever.

OpenStudy (jhannybean):

Were you able to solve this one?

OpenStudy (anonymous):

a/sin 54.9=160/sin109.3 .94a=(82)(160) a=140 140/sin54.9=c/sin15.8 (140)(.27)/.82 =.82c c=46 the answer is supposed to be 53 meters. My answers are no where near, I don't know what I have done incorrectly

OpenStudy (jhannybean):

a is already given.

OpenStudy (jhannybean):

Let's check.

OpenStudy (danjs):

i got 53.24

OpenStudy (danjs):

160/sin(180-15.8-109.3) = AB/sin(15.8)

OpenStudy (jhannybean):

\[\frac{160}{\sin(54.9)}=\frac{c}{\sin(15.8)}\]

OpenStudy (jhannybean):

\[160\sin(15.8) = c\sin(54.9)\]\[c=\frac{160\sin(15.8)}{\sin(54.9)} \approx 53.24\]

OpenStudy (jhannybean):

Rounding up to the nearest meter,you would have \(53~\text{meters}\)

OpenStudy (anonymous):

ok I was brain farting bad. using sub.. ok thank you very much. cheers

OpenStudy (jhannybean):

No problem :)